# A partition cover approach to tokenization

### v2

@@ Abstract
Tokenization is the process of encoding strings into tokens of a fixed vocabulary size, and is widely utilized in Natural Language Processing applications. The leading tokenization algorithm today is Byte-Pair Encoding (BPE), which formulates the tokenization problem as a compression problem and tackles it by performing sequences of merges. In this work, we formulate tokenization as an optimization objective, show that it is NP-hard via a simple reduction from vertex cover, and propose a polynomial-time greedy algorithm GreedTok. Our formulation naturally relaxes to the well-studied weighted maximum coverage problem which has a simple $(1-1/e)$-approximation algorithm GreedWMC. Through empirical evaluations on real-world corpora, we show that GreedTok outperforms BPE and Unigram on compression and achieves a covering score comparable to GreedWMC. Finally, our extensive pre-training for two transformer-based language models with 1 billion parameters, comparing the choices of BPE and GreedTok as the tokenizer, shows that GreedTok achieves a lower bit per byte even when we control for either the total dataset proportion or total training tokens.
@@ 1 Introduction
Tokenization encodes text into tokens from a fixed vocabulary and is fundamental to Natural Language Processing (NLP) applications. With Large Language Models (LLMs) growing in prominence, understanding tokenization has become increasingly important, as it plays an integral role in their architectures and even modest gains in efficiency can yield substantial computational savings. Furthermore, LLMs such as LLaMA [TLI+23] and Mistral [JSM+23] use fixed-length context windows, which will benefit from tokenizers with better compression utility that can fit more information within their context window. There are also prompt-based [WWS+22, YYZ+23] and fine-tuning [FTL+23] techniques to increase the number of processed tokens to improve model performance.
@@
A common way to formalize the tokenization problem is as a compression task that minimizes the ratio of tokens produced when tokenizing the input data. The leading tokenization algorithm today is Byte-Pair Encoding (BPE) [Gag94, SHB16, KHM+23], which formulates the tokenization problem as a compression problem and tackles it by performing a sequence of pairwise merges. Due to its popularity, there have been a multitude of recent works analyzing the theoretical properties of BPE [ZMG+23, KV24, WBP24]. Another approach frames tokenization as a pathing/sequence problem [Kud18, SRZ+24], Unigram [Kud18] is one such example and is favored by models with bidirectional contexts, i.e. [YDY+19] and [RSR+20]. Although Unigram is believed to perform better in pre-training tasks [BD20], its compression performance is weaker compared to BPE [SRZ+24], affecting its adoption for large-scale expensive pre-training in language modeling.
Contributions.
@@
In this work, we deviate from the usual path, sequence, and merge-based tokenization formulations. Instead, we propose to examine the tokenization problem as a problem of covering.
@@
C1. Partition cover formulation. We introduce a partition cover optimization formulation of the tokenization problem that goes beyond prior merge-based approaches [ZMG+23, KV24]. These earlier methods rely on bottom-up pairwise merges from an existing token set. However, the idea of tokenization is simply to efficiently represent a corpus with a judiciously selected set of tokens, whose construction is independent of such merge patterns. Our formulation subsumes these prior formulations in the following sense: all merge-based solutions are valid solutions to our formulation, but a solution to our formulation need not be based on bottom-up pairwise merges.
@@
C2. Simple NP-hardness proof. We provide a simple and intuitive proof that tokenization is NP-hard in our optimization objective formulation via a reduction from vertex cover [Kar72]. Although there has been a recent concurrent work [WBP24] that also showed that the tokenization problem is NP-hard, our proof is arguably simpler due to our formulation.
@@
C3. New polynomial-time tokenizer. We propose a polynomial-time greedy algorithm GreedTok that does not rely on a sequence of pairwise token merges or path construction. Evaluation on four real-world corpora shows that GreedTok outperforms BPE and Unigram with a stronger compression of $\sim$3tokens per word. Our implementation and compression evaluations can be found at https://github.com/PreferredAI/pcatt.
@@
C5. Empirical approximation of objective. Our partition cover formulation naturally relaxes to the well-studied weighted maximum coverage problem [Kar72, CK08] which has a simple $(1-1/e)$-approximation algorithm GreedWMC [Hoc96]. We empirically show that GreedTok and GreedWMC achieve a comparable objective function value for large $k$, despite the latter being a relaxed problem. Although a formal approximation guarantee for GreedTok currently escapes us, our analysis holds for practical scenarios encountered by NLP practitioners and this empirical method of investigating approximation guarantees may be of independent interest.
Related work.
There has been a recent surge of interest in analyzing tokenization. Zouhar et al. [ZMG+23] initiated a formal study of BPE using a bottom-up tokenization problem formulation that restricts token construction to sequential merges of two tokens from the existing vocubulary. Kozma and Voderholzer [KV24] proved that this bottom-up tokenization problem, and its more general variant, [^1] Referred to as optimal merge sequence and optimal pair encoding respectively in their work. is APX-complete using linear reduction [PY88] from maximum cut [Kar72] in cubic graphs. They also showed that BPE approximates a worst-case factor of between $0.333$ and $0.625$ for their general variant. In a recent concurrent work, Whittington, Bachmann, and Pimentel [WBP24] showed that both the bottom-up tokenization problem formulation and our partition cover formulation are NP-complete from the reduction of the maximum 2-satisfiability problem. Beyond theory, empirical studies such as [LBM23] and [SFWN23] have examined the practical downstream impact of tokenizer selection in NLP tasks. With regard to tokenizer-free architectures [TTR+22, YSF+23, PPR+24], our formulation draws the link to the fundamental binary decision problem of whether to merge adjacent characters, as one can view next-byte predictions as merging decisions.
Compared to prior and concurrent works, our NP-hardness proof (Theorem 1) is arguably simpler due to our formulation, and we contribute a novel tokenizer that is competitive in real-world scenarios.
Outline of paper.
We give our partition cover optimization formulation in Section 2, we prove that it is NP-hard in Section 3. GreedTok is designed in Section 4 and Section 5 contains our empirical evaluation against real world corpora. Finally, we conclude with some discussions in Section 6.
Notation.
[…]
@@ 2 A general optimization formulation for the tokenization problem
Let $\mathbf{\Sigma}$ be a fixed alphabet. A corpus $\mathcal{C}=(\mathbf{W},\textsc{count})$ consists of a set of words $\mathbf{W}\subseteq\mathbf{\Sigma}^{+}$ and a count function $\textsc{count}:\mathbf{W}\to\mathbb{N}$ that indicates word frequencies. Given a word $W\in\mathbf{\Sigma}^{+}$ and a set of tokens $\mathbf{S}\subseteq\mathbf{\Sigma}^{+}$, let $\textsc{partition}(W,\mathbf{S})$ denote the minimum number of tokens from $\mathbf{S}$ that can be concatenated in sequence to form $W$. For example, the word $W=\text{abc}$ can be tokenized as abc but not as acb, despite both token sets covering the same characters. The goal of minimizing the total number of tokens used to represent a corpus is often referred to as optimizing compression utility.
@@
Let $\mathbf{\Sigma}$ be a fixed alphabet and define the base token set $\mathbf{B}=\{(w):w\in\mathbf{\Sigma}\}$ as all singleton characters. Given a corpus $\mathcal{C}=(\mathbf{W},\textsc{count})$, a token budget $k\in\mathbb{N}_{>0}$, and a set of candidate tokens $\mathbf{T}\subseteq\mathbf{\Sigma}^{+}$, the goal is to find a subset $\mathbf{S}\subseteq\mathbf{T}$ such that $|\mathbf{S}|\leq k$ and $\sum_{W\in\mathbf{W}}\textsc{partition}(W,\mathbf{S}\cup\mathbf{B})\cdot\textsc{count}(W)$ is minimized.
@@
With the same setup as Problem 1, and given an additional integer threshold $\ell\in\mathbb{N}_{>0}$, determine whether there exists a subset $\mathbf{S}\subseteq\mathbf{T}$ such that $|\mathbf{S}|\leq k$ and $\sum_{W}^{\mathbf{W}}\textsc{partition}(W,\\ \mathbf{S}\cup\mathbf{B})\cdot\textsc{count}(W)\leq\ell$.
Our formulation differs subtly but significantly from prior tokenization models, such as those in [ZMG+23, KV24]. Rather than viewing tokenization through the lens of string compression algorithms, we reduce tokenization to the following question: Are two adjacent singleton symbols within a string covered by the same token? This perspective emphasizes that tokenization is fundamentally about selecting a compact vocabulary that fully covers the corpus without imposing algorithmic constraints such as bottom-up merge sequences, which are artifacts of specific approaches like BPE. We further elaborate on the implications of our formulation in Section 4.2.
@@ 3 Tokenization is NP-hard
In this section, we prove that the tokenization decision problem (Problem 2) is NP-hard by a reduction from the vertex cover problem, which is known to be NP-hard [Kar72]. Given a graph $\mathcal{G}=(\mathbf{V},\mathbf{E})$ with vertex set $\mathbf{V}$ and edge set $\mathbf{E}$, a vertex cover is a subset $\mathbf{S}\subseteq\mathbf{V}$ of vertices such that $|\mathbf{S}\cap\{U,V\}|\geq 1$ for every edge $\{U,V\}\in\mathbf{E}$. The decision variant of the vertex cover problem asks whether there exists a subset $\mathbf{S}$ of size at most $k$.
@@
We will prove this by reducing the vertex cover problem, which is known to be NP-hard [Kar72], to the tokenization problem. Given an arbitrary vertex cover problem instance, we show how to construct a corresponding tokenization instance. Then, we argue that the derived tokenization problem instance is a YES instance if and only if the original vertex cover problem instance is a YES instance. In this proof, for clarity, we will write words $W\in\mathbf{W}$ as a tuple of singletons instead of usual plaintext, e.g. (h,e,l,l,o) instead of hello.
Construction.
@@
Consider an arbitrary vertex cover problem given by the graph $\mathcal{G}=(\mathbf{V},\mathbf{E})$ over $n$ vertices $\mathbf{V}=\{V_{1},\ldots,V_{n}\}$ and a positive integer $k\in\mathbb{N}_{\geq 0}$. To construct an instance of the tokenization problem, we first define the alphabet as follows: $\mathbf{\Sigma}=\{V_{1},\ldots,V_{n},@\}$ where @@@@ is an additional symbol which we will use later. So, we have $\mathbf{B}=\{(V_{1}),\ldots,(V_{n}),(@)\}$. For each edge $\{V_{i},V_{j}\}\in\mathbf{E}$ with $i<j$, we create a word $W_{i,j}=(@,V_{i},@,V_{j},@)$ and define the set of words as $\mathbf{W}=\{W_{i,j}:\{V_{i},V_{j}\}\in\mathbf{E}\}$ where each word has count 1, i.e. $\textsc{count}(W)=1$ for all $W\in\mathbf{W}$. Then, we define the set of candidate tokens $\mathbf{T}=\{(@,V_{i},@):V_{i}\in\mathbf{V}\}$. Finally, we set $\ell=3|\mathbf{W}|=3|\mathbf{E}|$ and associate the parameter $k$ in the vertex cover problem instance to the corresponding parameter $k$ in the tokenization problem instance. One can check that this derived tokenization instance can be constructed in polynomial time.
Observation.
@@
Observe that every word $W\in\mathbf{W}$ has length 5 and each token in $\mathbf{S}$ has length 3, so $\textsc{partition}(W,\mathbf{S}\cup\mathbf{B})$ will either be 3, when there is some token in $\mathbf{S}$ that is a contiguous subword of $W$, or 5 otherwise. For instance, given the word $W_{i,j}=(@,V_{i},@,V_{j},@)$, we have $\textsc{partition}(W_{i,j},\mathbf{S}\cup\mathbf{B})=3$ if and only if at least one of (@$,V_{i},@)$ or (@$,V_{j},@)$ is chosen in $\mathbf{S}$ (both could be in $\mathbf{S}$). Furthermore, since all words have count 1, the tokenization problem becomes finding $\mathbf{S}\subseteq\mathbf{T}$ such that $|\mathbf{S}|\leq k$ and $\sum_{W\in\mathbf{W}}\textsc{partition}(W,\mathbf{S}\cup\mathbf{B})\leq\ell=3|\mathbf{W}|$.
YES instance of tokenization problem to YES instance of vertex cover.
@@
Suppose there exists a subset $\mathbf{S}\subseteq\mathbf{T}$ of tokens such that $|\mathbf{S}|\leq k$ and $\sum_{W\in\mathbf{W}}\textsc{partition}(W,\mathbf{S}\cup\mathbf{B})\leq\ell=3|\mathbf{W}|$. Then, from the observation above, we know that this can only happen when $\textsc{partition}(W,\mathbf{S}\cup\mathbf{B})=3$ for every $W\in\mathbf{W}$. This implies that for each word $W_{i,j}$, at least one of (@$,V_{i},@)$ or (@$,V_{j},@)$ is chosen in $\mathbf{S}$. Therefore, $\mathbf{S}_{\mathcal{G}}=\{V_{i}\in\mathbf{V}:(@,V_{i},@)\in\mathbf{S}\}$ is a subset of size $|\mathbf{S}_{\mathcal{G}}|=|\mathbf{S}|\leq k$ and corresponds to a vertex cover of the original graph $\mathcal{G}$.
YES instance of vertex cover to YES instance of tokenization problem.
@@
Suppose the original vertex cover instance for graph $\mathcal{G}=(\mathbf{V},\mathbf{E})$ has a vertex cover $\mathbf{S}_{\mathcal{G}}$ of size $|\mathbf{S}_{\mathcal{G}}|\leq k$. Then, let us define $\mathbf{S}=\{(@,V_{i},@)\in\mathbf{\Sigma}^{+}:V_{i}\in\mathbf{S}_{\mathcal{G}}\}$ as the set of chosen tokens of size $|\mathbf{S}|=|\mathbf{S}_{\mathcal{G}}|\leq k$. Since $\mathbf{S}_{\mathcal{G}}$ is a set cover for $\mathcal{G}$, by construction of $\mathbf{W}$, we see that $\textsc{partition}(W,\mathbf{S}\cup\mathbf{B})=3$ for all $W\in\mathbf{W}$. Therefore, $\sum_{W\in\mathbf{W}}\textsc{partition}(W,\mathbf{S}\cup\mathbf{B})=3|\mathbf{W}|$. ∎

### v3

# A Partition Cover Approach for Tokenization

@@ Abstract
Tokenization is the process of encoding strings into tokens of a fixed vocabulary size, and is widely utilized in Natural Language Processing applications. The leading tokenization algorithm today is Byte-Pair Encoding (BPE), which formulates the tokenization problem as a compression problem and tackles it by performing sequences of merges. In this work, we formulate tokenization as an optimization objective, show that it is NP-hard via a simple reduction from vertex cover, and propose a polynomial-time greedy algorithm GreedTok. Our formulation naturally relaxes to the well-studied weighted maximum coverage problem which has a simple (11/e)(1-1/e)-approximation algorithm GreedWMC. Through empirical evaluations on real-world corpora, we show that GreedTok outperforms BPE and Unigram on compression and achieves a covering score comparable to GreedWMC. Finally, our extensive pre-training for two transformer-based language models with 1 billion parameters, comparing the choices of BPE and GreedTok as the tokenizer, shows that GreedTok achieves a lower bit per byte even when we control for either the total dataset proportion or total training tokens.
@@ 1 Introduction
Tokenization encodes text into tokens from a fixed vocabulary and is fundamental to Natural Language Processing (NLP) applications. With Large Language Models (LLMs) growing in prominence, understanding tokenization has become increasingly important, as it plays an integral role in their architectures and even modest gains in efficiency can yield substantial computational savings. Furthermore, LLMs such as LLaMA TLI+ (23) and Mistral JSM+ (23) use fixed-length context windows, which will benefit from tokenizers with better compression utility that enables them to fit more information within their context window. There are also prompt-based WWS+ (22); YYZ+ (23) and fine-tuning FTL+ (23) techniques that increase the number of processed tokens to improve model performance.
@@
A common way to formalize the tokenization problem is as a compression task that minimizes the ratio of tokens produced when tokenizing the input data. The leading tokenization algorithm today is Byte-Pair Encoding (BPE) Gag (94); SHB (16); KHM+ (23), which formulates the tokenization problem as a compression problem and tackles it by performing a sequence of pairwise merges. Due to its popularity, there have been a multitude of recent works analyzing the theoretical properties of BPE ZMG+ (23); KV (24); WBP (24). Another approach frames tokenization as a pathing/sequence problem Kud (18); SRZ+ (24), Unigram Kud (18) is one such example and is favored by models with bidirectional contexts, i.e. YDY+ (19) and RSR+ (20). Although Unigram is believed to perform better in pre-training tasks BD (20), its compression performance is weaker compared to BPE SRZ+ (24), affecting its adoption for large-scale expensive pre-training in language modeling.
@@
Contributions. In this work, we deviate from the usual path, sequence, and merge-based tokenization formulations. Instead, we propose to examine the tokenization problem as a problem of covering.
@@
C1. Partition cover formulation. We introduce a partition cover optimization formulation of the tokenization problem that goes beyond prior merge-based approaches ZMG+ (23); KV (24). These earlier methods rely on bottom-up pairwise merges from an existing token set. However, the idea of tokenization is simply to efficiently represent a corpus with a judiciously selected set of tokens, whose construction is independent of such merge patterns. Our formulation subsumes these prior formulations in the following sense: all merge-based solutions are valid solutions to our formulation, but a solution to our formulation need not be based on bottom-up pairwise merges.
@@
C2. Simple NP-hardness proof. We provide a simple and intuitive proof that tokenization is NP-hard in our optimization objective formulation via a reduction from vertex cover Kar (72). Although there has been a recent concurrent work WBP (24) that also showed that the tokenization problem is NP-hard, our proof is arguably simpler due to our formulation.
@@
C3. New polynomial-time tokenizer. We propose a polynomial-time greedy algorithm GreedTok that does not rely on a sequence of pairwise token merges or path construction. Evaluation on four real-world corpora shows that GreedTok outperforms BPE and Unigram with a stronger compression of $\sim$3% tokens per word. Our implementation and compression evaluations can be found at https://github.com/PreferredAI/pcatt/; see supplementary materials.
@@
C5. Empirical approximation of objective. Our partition cover formulation naturally relaxes to the well-studied weighted maximum coverage problem Kar (72); CK (08) which has a simple (11/e)(1-1/e)-approximation algorithm GreedWMC Hoc (96). We empirically show that GreedTok and GreedWMC achieve a comparable objective function value for large $k$, despite the latter being a relaxed problem. Although a formal approximation guarantee for GreedTok currently escapes us, our analysis holds for practical scenarios encountered by NLP practitioners and this empirical method of investigating approximation guarantees may be of independent interest.
Related work. There has been a recent surge of interest in analyzing tokenization. ZMG+ (23) initiated a formal study of BPE using a bottom-up tokenization problem formulation that restricts token construction to sequential merges of two tokens from the existing vocubulary. KV (24) proved that this bottom-up tokenization problem, and its more general variant, [^1] Referred to as optimal merge sequence and optimal pair encoding respectively in their work. is APX-complete using linear reduction PY (88) from maximum cut Kar (72) in cubic graphs. They also showed that BPE approximates a worst-case factor of between $0.333$ and $0.625$ for their general variant. In a recent concurrent work, WBP (24) showed that both the bottom-up tokenization problem formulation and our partition cover formulation are NP-complete from the reduction of the maximum 2-satisfiability problem. Beyond theory, empirical studies such as LBM (23) and SFWN (23) have examined the practical downstream impact of tokenizer selection in NLP tasks. With regard to tokenizer-free architectures TTR+ (22); YSF+ (23); PPR+ (24), our formulation draws the link to the fundamental binary decision problem of whether to merge adjacent characters, as one can view next-byte predictions as merging decisions.
Compared to prior and concurrent works, our NP-hardness proof (Theorem˜1) is arguably simpler due to our formulation, and we contribute a novel tokenizer that is competitive in real-world scenarios.
Outline of paper. We give our partition cover optimization formulation in Section˜2, we prove that it is NP-hard in Section˜3. GreedTok is designed in Section˜4 and Section˜5 contains our empirical evaluation against real world corpora. Finally, we conclude with some discussions in Section˜6.
Notation. We use standard set notations such as $|\mathbf{A}|$ to represent the size of set $\mathbf{A}$, and standard asymptotic notations such as $O(\cdot)$. Numbers are represented with small letters, strings/words with capital letters, and sets with bold letters. Unordered sets are denoted by $\{\cdot\}$ and ordered tuples are denoted by $(\cdot)$. We describe words in plaintext, e.g., hello, or as a tuple of singletons, e.g., (h,e,l,l,o).
@@ 2 A general optimization formulation for the tokenization problem
Let $\mathbf{\Sigma}$ be a fixed alphabet , i.e., a basic character set. This may be the 26 lowercase English letters or the full Unicode set, depending on the context. A corpus $\mathcal{C}=(\mathbf{W},\textsc{count})$ consists of a set of distinct words $\mathbf{W}\subseteq\mathbf{\Sigma}^{+}$, and a count function $\textsc{count}:\mathbf{W}\to\mathbb{N}$ that indicates word frequencies. Given a word $W\in\mathbf{\Sigma}^{+}$ and a set of tokens $\mathbf{S}\subseteq\mathbf{\Sigma}^{+}$, let $\textsc{partition}(W,\mathbf{S})$ denote the minimum number of tokens from $\mathbf{S}$ that can be concatenated in sequence to form $W$. For example, the word $W=\text{abc}$ can be tokenized as abc but not as acb, despite both token sets covering the same characters.
The goal of minimizing the total number of tokens used to represent a corpus is often referred to as optimizing compression utility.
@@
Let $\mathbf{\Sigma}$ be a fixed alphabet and define the base token set $\mathbf{B}=\{(w):w\in\mathbf{\Sigma}\}$ as all singleton characters. Given a corpus $\mathcal{C}=(\mathbf{W},\textsc{count})$, a token budget $k\in\mathbb{N}_{>0}$, and a set of candidate tokens $\mathbf{T}\subseteq\mathbf{\Sigma}^{+}$, the goal is to find a subset $\mathbf{S}\subseteq\mathbf{T}$ such that $|\mathbf{S}|\leq k$ and $\sum_{W\in\mathbf{W}}\textsc{partition}(W,\mathbf{S}\cup\mathbf{B})\cdot\textsc{count}(W)$ is minimized. Note that tokens $\mathbf{T}$ are candidate substrings drawn from $\mathbf{\Sigma}^{*}$ that may or may not correspond to actual words.
@@
With the same setup as Problem˜1, and given an additional integer threshold $\ell\in\mathbb{N}_{>0}$, determine whether there exists a subset $\mathbf{S}\subseteq\mathbf{T}$ such that $|\mathbf{S}|\leq k$ and $\sum_{W\in\mathbf{W}}\textsc{partition}(W,\mathbf{S}\cup\mathbf{B})\cdot\textsc{count}(W)\leq\ell$.
Our formulation differs subtly but significantly from prior tokenization models, such as those in ZMG+ (23); KV (24). Rather than viewing tokenization through the lens of string compression algorithms, we reduce tokenization to the following question: Are two adjacent singleton symbols within a string covered by the same token? This perspective emphasizes that tokenization is fundamentally about selecting a compact vocabulary that fully covers the corpus without imposing algorithmic constraints such as bottom-up merge sequences, which are artifacts of specific approaches like BPE. We further elaborate on the implications of our formulation in Section˜4.1.
@@ 3 Tokenization is NP-hard
In this section, we prove that the tokenization decision problem (Problem˜2) is NP-hard by a reduction from the vertex cover problem, which is known to be NP-hard Kar (72). Given a graph $\mathcal{G}=(\mathbf{V},\mathbf{E})$ with vertex set $\mathbf{V}$ and edge set $\mathbf{E}$, a vertex cover is a subset $\mathbf{S}\subseteq\mathbf{V}$ of vertices such that $|\mathbf{S}\cap\{U,V\}|\geq 1$ for every edge $\{U,V\}\in\mathbf{E}$. The decision variant of the vertex cover problem asks whether there exists a subset $\mathbf{S}$ of size at most $k$.
@@
We will prove this by reducing the vertex cover problem, which is known to be NP-hard Kar (72), to the tokenization problem. Given an arbitrary vertex cover problem instance, we show how to construct a corresponding tokenization instance. Then, we argue that the derived tokenization problem instance is a YES instance if and only if the original vertex cover problem instance is a YES instance. In this proof, for clarity, we will write words $W\in\mathbf{W}$ as a tuple of singletons instead of usual plaintext, e.g. (h,e,l,l,o) instead of hello.
@@
Construction. Consider an arbitrary vertex cover problem given by the graph $\mathcal{G}=(\mathbf{V},\mathbf{E})$ over $n$ vertices $\mathbf{V}=\{V_{1},\ldots,V_{n}\}$ and a positive integer $k\in\mathbb{N}_{\geq 0}$. To construct an instance of the tokenization problem, we first define the alphabet as follows: $\mathbf{\Sigma}=\{V_{1},\ldots,V_{n},@\}$ where @@ is an additional symbol which we will use later. So, we have $\mathbf{B}=\{(V_{1}),\ldots,(V_{n}),(@)\}$. For each edge $\{V_{i},V_{j}\}\in\mathbf{E}$ with $i<j$, we create a word $W_{i,j}=(@,V_{i},@,V_{j},@)$ and define the set of words as $\mathbf{W}=\{W_{i,j}:\{V_{i},V_{j}\}\in\mathbf{E}\}$ where each word has count 1, i.e. $\textsc{count}(W)=1$ for all $W\in\mathbf{W}$. Then, we define the set of candidate tokens $\mathbf{T}=\{(@,V_{i},@):V_{i}\in\mathbf{V}\}$. Finally, we set $\ell=3|\mathbf{W}|=3|\mathbf{E}|$ and associate the parameter $k$ in the vertex cover problem instance to the corresponding parameter $k$ in the tokenization problem instance. One can check that this derived tokenization instance can be constructed in polynomial time.
@@
Observation. Observe that every word $W\in\mathbf{W}$ has length 5 and each token in $\mathbf{S}$ has length 3, so $\textsc{partition}(W,\mathbf{S}\cup\mathbf{B})$ will either be 3, when there is some token in $\mathbf{S}$ that is a contiguous subword of $W$, or 5 otherwise. For instance, given the word $W_{i,j}=(@,V_{i},@,V_{j},@)$, we have $\textsc{partition}(W_{i,j},\mathbf{S}\cup\mathbf{B})=3$ if and only if at least one of $(@,V_{i},@)$ or $(@,V_{j},@)$ is chosen in $\mathbf{S}$ (both could be in $\mathbf{S}$). Furthermore, since all words have count 1, the tokenization problem becomes finding $\mathbf{S}\subseteq\mathbf{T}$ such that $|\mathbf{S}|\leq k$ and $\sum_{W\in\mathbf{W}}\textsc{partition}(W,\mathbf{S}\cup\mathbf{B})\leq\ell=3|\mathbf{W}|$.
@@
YES instance of tokenization problem to YES instance of vertex cover. Suppose there exists a subset $\mathbf{S}\subseteq\mathbf{T}$ of tokens such that $|\mathbf{S}|\leq k$ and $\sum_{W\in\mathbf{W}}\textsc{partition}(W,\mathbf{S}\cup\mathbf{B})\leq\ell=3|\mathbf{W}|$. Then, from the observation above, we know that this can only happen when $\textsc{partition}(W,\mathbf{S}\cup\mathbf{B})=3$ for every $W\in\mathbf{W}$. This implies that for each word $W_{i,j}$, at least one of $(@,V_{i},@)$ or $(@,V_{j},@)$ is chosen in $\mathbf{S}$. Therefore, $\mathbf{S}_{\mathcal{G}}=\{V_{i}\in\mathbf{V}:(@,V_{i},@)\in\mathbf{S}\}$ is a subset of size $|\mathbf{S}_{\mathcal{G}}|=|\mathbf{S}|\leq k$ and corresponds to a vertex cover of the original graph $\mathcal{G}$.
@@
YES instance of vertex cover to YES instance of tokenization problem. Suppose the original vertex cover instance for graph $\mathcal{G}=(\mathbf{V},\mathbf{E})$ has a vertex cover $\mathbf{S}_{\mathcal{G}}$ of size $|\mathbf{S}_{\mathcal{G}}|\leq k$. Then, let us define $\mathbf{S}=\{(@,V_{i},@)\in\mathbf{\Sigma}^{+}:V_{i}\in\mathbf{S}_{\mathcal{G}}\}$ as the set of chosen tokens of size $|\mathbf{S}|=|\mathbf{S}_{\mathcal{G}}|\leq k$. Since $\mathbf{S}_{\mathcal{G}}$ is a set cover for $\mathcal{G}$, by construction of $\mathbf{W}$, we see that $\textsc{partition}(W,\mathbf{S}\cup\mathbf{B})=3$ for all $W\in\mathbf{W}$. Therefore, $\sum_{W\in\mathbf{W}}\textsc{partition}(W,\mathbf{S}\cup\mathbf{B})=3|\mathbf{W}|$. ∎